OpenStudy (anonymous):

How would you do the reciprocal of 5/2 over 1?

OpenStudy (anonymous):

5/2 / 1 = 5/2 Reciprocal is just one over that: 1 / (5 / 2) = 1 / 5 / 2 = 2 / 5